# You Only Acquire Sparse-channel (YOAS): A Unified Framework for Dense-channel EEG Generation

### v1

@@ III-C Data Preprocessing
Given the low signal-to-noise ratio (SNR) of EEG signals and the potential for signal corruption, it is likely that the bias signal data $\{\beta_{ij}\}_{i=1,j=2}^{K,\hat{M}}$ derived from the Data Preparation stage of Fig. 1 may contain erroneous data. To mitigate these challenges, the following preprocessing procedures have been implemented. Initially, outlier removal is conducted by empirically eliminating data points within each signal segment that deviate by one or two standard deviations from the total segment signal of each channel. Subsequently, linear interpolation is applied to address missing data points. Additionally, NaN values in $\{\beta_{ij}\}_{i=1,j=2}^{K,\hat{M}}$ are filled using linear interpolation, and any remaining INF values post-interpolation are replaced with zero. Finally, Multiscale Principal Components Analysis (MSPCA) [23] is employed to reduce dimensionality and extract effective features from the interpolated data.
@@ III-D Biased-EEG Generation
Figure 2: Illustration of hypotheses for multi-channel EEG generation pathways: (A) Direct one-way generation involves generating signals directly with reference to the channel; (B) Indirect one-way generation requires an intermediate channel for the reference channel to generate signals indirectly; (C) Mutual generation involves two channels serving as reference channels for each other, facilitating mutual signal generation.
Figure 3: BiasEEGGanFormer employs an adversarial generation architecture, where the Biased-EEGGenerator generates data fed into the Biased-EEGDiscriminator for loss feedback, thereby improving the generation process quality. The resulting output comprises the One-stage Generated Bias, enhancing EEG signal refinement.
Figure 4: BiasEEGDiffFormer aligns One-stage Generated Bias with the correspondingly physical locations, first applying one-dimensional convolution in ConvBlock for feature extraction, and subsequently involving DiffusionBlock to generate Two-stage Generated Bias based on Regional Reference Signal and a diffusion model.
@@
1) BiasEEGGanFormer
Compared with the WGAN series [30], the Transformer model [24] adopts a self-attention mechanism, enabling it to capture long-term dependencies in time series data without being constrained by sequence length. Meanwhile, compared with traditional GAN models, the Transformer architecture mitigates issues such as generator collapse and training instability [30]. To generate quasi-realistic data, we propose BiasEEGGanFormer as shown in Fig. 3, which retains the original GAN architecture while incorporating the Transformer. Specifically, inspired by [31], we opt for a simpler approach, adding linear and ReLU layers after the Transformer structure. For the generator, an embedding step extracts features from input data, followed by the Multi-Head Self-Attention module capturing sequence dependencies. Subsequently, an MLP with two linear layers and ReLU activation further processes encoded and decoded outputs, with data generated through a fully connected layer. Both generator and discriminator share the same structure for balance, producing data termed as ’One-stage Generated Bias’ i.e. $b_{ij}$ described in Formula 6. Unlike previous works [30],[32],[33] employing numerous convolutional layers or complex up-sampling and down-sampling techniques, our BiasEEGGanFormer prioritizes the generation effectiveness and efficiency.
2) BiasEEGDiffFormer
Taking advantage of the correlation between EEG signals, electrode positions, and reference channel signals, BiasEEGDiffFormer (Fig. 4) is developed as an extension of the BiasEEGGanFormer. Initially, we align each One-stage Generated Bias $b_{ij}$ with the reference signal $O_{i1}$ of its associated physical location $RD_{i}$. By merging these generated data with reference signals, we construct a virtual two-dimensional EEG distribution map, which enhances the interpretability of generated data $C_{ij}$ for each channel. Next, we utilize the ConvBlock to further extract features, followed by DiffusionBlock [28]. Drawing upon the regional reference signal, we meticulously select the generated data that most accurately aligned with its position, resulting in ’Two-stage Generated Bias’ i.e. $B_{ij}$ describled in Formula 7 with a more detailed expression as
@@
by repeatedly applying the backward denoising process, where the final signal $B_{ij}$ is gradually generated from the noise. Thus, the generation process of $B_{i,j}$ can be rewritten to Algorithm 1 as below.
Algorithm 1 Generation Process of $B_{ij}$
@@
Num layers 6
Num heads 8
III-E Synthetic EEG Generation
As shown in Fig. 1, to pursue optimized sets of generation paths for generating multi-channel EEG signals across the whole brain, we propose a deduction paradigm consisting of two consecutive paradigms: i.e, Synthetic Paradigm 1 (referring to Algorithm 2) and Synthetic Paradigm 2 (referring to Algorithm 3). Given the initial regional division’s insufficient accuracy in Section III-B Data Preparation, Synthetic Paradigm 1 aims to refine signal generation paths, including direct, indirect, and mutual generation among channels. Building upon the refined results of Synthetic Paradigm 1, Synthetic Paradigm 2 is designed to retrieve the final optimized sets of the generation paths and regional divisions for generating dense-channel EEG signals across the whole brain.
IV EXPERIMENTS
IV-A Implementation Details
Datasets: We conduct training and validation of our models on the FACED dataset provided in reference [34]. The dataset comprises recordings from 123 subjects, each with 32 EEG channels, while they watched 28 video clips, all lasting for 30 seconds. These clips cover a range of emotions, where we select four clips for neutral emotion. In the experiments, we select 250Hz as the sampling frequency. There are a total of 55 subjects, including 30 subjects for training, and 25 subjects for testing.
Training Details: We conduct training and testing on the provided dataset using the same parameters outlined in Table LABEL:tab:table1. Initially, for the data utilized during training, we employ WGAN-GP [30] to assess the feasibility of each generation path showed in Fig. 5 through the testing of one-to-one channel signal generation. Subsequently, compared with WGAN-GP, we retest the performance of the paths generated by our BiasEEGGanFormer. All computations are executed on NVIDIA Tesla V100, and the Early-Stopping strategy is applied in the training stage.
@@ IV-B Performance Comparison
Figure 6: Performance comparison between BiasEEGGanFormer (orange) and WGAN-GP (blue): (A) and (B) demonstrate that BiasEEGGanFormer takes only 22.22% of the time cost of WGAN-GP, while maintaining comparable generation quality. (C) and (D) provide further evidence, showing that BiasEEGGanFormer outperforms WGAN-GP in quality across 41 generation path experiments, with less time cost.
Figure 7: The power spectrum comparison between the “indifferent generated signal” of WGAN-GP in (B), and BiasEEGGanformer in (C), demonstrates BiasEEGGanformer’s superior performance in attenuating high-amplitude components in frequency-spectrum domain, potentially indicative of noise.
@@
IV-C Optimized Generation Paths
Based on the final output of the YOAS framework in our experiments, we observed that the initial ten regional divisions depicted in Fig. 5 were optimized and merged into seven regional divisions on neutral emotion data, as demonstrated in Fig. 8 following Synthetic Paradigm 2. Specifically, as illustrated in Fig. 8(A), Fp1 can directly generate F3, F7, FC5, and Fp2, while FP2 indirectly generates F4, F8, and FC6 through Fp1. Notably, F7 can directly generate FC5, and F8 can directly generate FC6. In Fig. 8(B), CP1 generates T3 exclusively, CP2 generates T4 exclusively, while pair of O1 and O2 reciprocally generate each other, as well as pair of A1 and A2, can generate each other mutually. Referring to Fig. 8(C), Fz directly generates FC1, FC2, and Cz, while indirectly generating C3 and C4 through Cz. As shown in Fig. 8(D), Pz directly generates Cp5, Cp6, P4, and PO4, with indirect generation of P3 through P4. Notably, PO4, PO3, and Oz mutually generate each other, with Oz also generating T5 and T6, both of which mutually generate each other.

### v2

@@ III-C Data Preprocessing
Given the low signal-to-noise ratio (SNR) of EEG signals and the potential for signal corruption, it is likely that the bias signal data $\{\beta_{ij}\}_{i=1,j=2}^{K,\hat{M}}$ derived from the Data Preparation stage of Fig. 1 may contain erroneous data. To mitigate these challenges, the following preprocessing procedures have been implemented. Initially, outlier removal is conducted by empirically eliminating data points within each signal segment that deviate by one or two standard deviations from the total segment signal of each channel. Subsequently, linear interpolation is applied to address missing data points. Additionally, NaN values in $\{\beta_{ij}\}_{i=1,j=2}^{K,\hat{M}}$ are filled using linear interpolation, and any remaining INF values are replaced with zero. Finally, Multiscale Principal Components Analysis (MSPCA) [37] is employed to reduce dimensionality and extract effective features from the interpolated data.
@@ III-D Biased-EEG Generation
Figure 2: Hypotheses illustration for dense-channel EEG generation pathways: (A) Direct one-way generation involves generating signals directly with reference to the channel; (B) Indirect one-way generation requires an intermediate channel for the reference channel to generate signals indirectly; (C) Mutual generation involves two channels serving as reference channel mutually.
Figure 3: BiasEEGGanFormer employs an adversarial generation architecture, where the Biased-EEGGenerator generates data fed into the Biased-EEGDiscriminator for loss feedback, thereby improving the generation quality. The resulting output comprises the One-stage Generated Bias, enhancing EEG signal refinement.
Figure 4: BiasEEGDiffFormer aligns One-stage Generated Bias with the correspondingly physical locations, first applying 1D convolution in ConvBlock for feature extraction, and subsequently involving DiffusionBlock to generate Two-stage Generated Bias based on Regional Reference Signal and diffusion model.
@@
BiasEEGGanFormer. Compared with the WGAN series [44], the Transformer model [38] adopts a self-attention mechanism, enabling it to capture long-term dependencies in time series data without being constrained by sequence length. Compared with traditional GAN models, the Transformer mitigates issues such as generator collapse and training instability [44]. To generate quasi-realistic data, we propose BiasEEGGanFormer as shown in Fig. 3, which retains the original GAN architecture while incorporating the Transformer. Specifically, inspired by [45], we opt for a simpler approach, adding linear and ReLU layers after the Transformer structure. For the generator, an embedding step extracts features from input data, followed by the Multi-Head Self-Attention module capturing sequence dependencies. Subsequently, an MLP with two linear layers and ReLU activation further processes encoded and decoded outputs, with data generated through a fully connected layer. Both generator and discriminator share the same structure for balance, producing data termed as ‘One-stage Generated Bias’ i.e. $b_{ij}$ described in Formula 6. Unlike previous works [44],[46],[47] employing numerous convolutional layers or complex up-sampling and down-sampling techniques, our BiasEEGGanFormer prioritizes the generation effectiveness and efficiency.
BiasEEGDiffFormer. Taking advantage of the correlation between EEG signals, electrode positions, and reference channel signals, BiasEEGDiffFormer (Fig. 4) is developed as an extension of BiasEEGGanFormer. Initially, we align each One-stage Generated Bias $b_{ij}$ with the reference signal $O_{i1}$ of its associated physical location $RD_{i}$. By merging these generated data with reference signals, we construct a virtual two-dimensional EEG distribution map, which enhances the interpretability of generated data $C_{ij}$ for each channel. Next, we utilize the ConvBlock to further extract features, followed by DiffusionBlock [42]. Drawing upon the regional reference signal, we meticulously select the generated data that most accurately aligned with its position, resulting in ‘Two-stage Generated Bias’ i.e. $B_{ij}$ describled in Formula 7 with a more detailed expression as
@@
by repeatedly applying the backward denoising process, where the final signal $B_{ij}$ is gradually generated from the noise. Thus, the generation process of $B_{i,j}$ can be rewritten to Algorithm 1.
Algorithm 1 : Generation Process of $B_{ij}$
@@
III-E Synthetic EEG Generation
As shown in Fig. 1, to pursue optimized generation paths for generating dense-channel EEG signals across the whole brain, we propose a deduction paradigm consisting of two consecutive paradigms: i.e, Synthetic Paradigm 1 (referring to Algorithm 2) and Synthetic Paradigm 2 (referring to Algorithm 3). Given the initial regional division’s insufficient accuracy in Section III-B Data Preparation, Synthetic Paradigm 1 aims to refine signal generation paths, including direct, indirect, and mutual generation among channels. Building upon the refined results of Synthetic Paradigm 1, Synthetic Paradigm 2 is designed to retrieve the final optimized sets of the generation paths and regional divisions for generating dense-channel EEG signals across the whole brain.
IV EXPERIMENTS
IV-A Implementation Details
Datasets: We conduct training and validation of our models on the FACED dataset provided in reference [48]. The dataset comprises recordings from 123 subjects, each with 32 EEG channels, while they watched 28 video clips, all lasting for 30 seconds. These clips cover nine emotions, including Anger, Disgust, Fear, Sadness, Neutral, Amusement, Inspiration, Joy and Tenderness. Each positive and negative emotion is represented by three clips, while the neutral emotion is depicted in four clips. In the experiments, we select 250Hz as the sampling frequency, resulting in a total of 55 subjects, including 30 subjects for training, and 25 subjects for testing.
Training Details: We conduct training and testing on the provided dataset using the same parameters outlined in Table LABEL:tab:table1. Initially, for the data utilized during training, we employ SpecGAN [49], WaveCycleGAN-GP [50] and WGAN-GP[44] respectively to assess the feasibility of each generation path showed in Fig. 5 through the testing of one-to-one channel signal generation. Subsequently, we retest the performance of the paths generated by our BiasEEGGanFormer. All computations are executed on NVIDIA Tesla V100, and the Early-Stopping strategy is applied in the training stage.
@@ IV-A Implementation Details
Layer Number 6
Head Number 8
@@ IV-B Performance Comparison
IV-C Optimized Generation Paths on Neutral Emotion Data
Based on the output of the YOAS framework in our experiments, we observed that the initial ten regional divisions depicted in Fig. 5 were optimized and merged into seven regional divisions on neutral emotion data, as demonstrated in Fig. 8 following Synthetic Paradigm 2. Specifically, as illustrated in Fig. 8(A), Fp1 can directly generate F3, F7, FC5, and Fp2, while FP2 indirectly generates F4, F8, and FC6 through Fp1. Notably, F7 can directly generate FC5, and F8 can directly generate FC6. In Fig. 8(B), CP1 generates T3 exclusively, CP2 generates T4 exclusively, while pair of O1 and O2 reciprocally generate each other, as well as pair of A1 and A2, can generate each other mutually. Referring to Fig. 8(C), Fz directly generates FC1, FC2, and Cz, while indirectly generating C3 and C4 through Cz. As shown in Fig. 8(D), Pz directly generates Cp5, Cp6, P4, and PO4, with indirect generation of P3 through P4. Notably, PO4, PO3, and Oz mutually generate each other, with Oz also generating T5 and T6, both of which mutually generate each other.
Figure 6: Performance comparison between BiasEEGGanFormer, SpecGAN [49], WaveCycleGAN-GP [50] and WGAN-GP[44]: (A) and (B) demonstrate that BiasEEGGanFormer is swifter to train than other models while maintaining comparable generation quality. (C) provides further evidence, showing that BiasEEGGanFormer outperforms other models in quality across 42 generation path experiments, with less time cost.
Figure 7: The power spectrum comparison between the “indifferent generated signal” of BiasEEGGanformer in (B), SpecGAN [49] in (C), WaveCycleGAN-GP[50] in (D), and WGAN-GP[44] in (E), demonstrates BiasEEGGanformer’s superior performance in attenuating high-amplitude components in frequency-spectrum domain, potentially indicative of noise.
@@ IV-C Optimized Generation Paths on Neutral Emotion Data
Figure 9: Multi-channel EEG signal generation paradigms on neutral, negative and positive emotion data. (I)-(IV) are the optimized generation paths for negative emotion, depicting generation paths of anger, disgust, fear, and sadness respectively. (V) is the optimized generation paths for neutral emotion. (VI)-(IX) are the optimized generation paths for positive emotion, portraying generation paths of amusement, inspiration, joy, and tenderness, in that order.
IV-D Optimized Generation Paths on Negative Emotion Data
In Fig. 9, (I)–(IV) are the negative emotion generation paradigms. It is worth noting that (I) and (II) share similar generation paradigms, as do (III) and (IV). In (A) of (I)-(IV) is mainly with Fp1 as reference channel, and F8 as reference channel in one-way generation paradigm alone, while in (A) of (III) and (IV), F8 can be derived by negating the F7 signal value generated from Fp1, which further reduces the number of reference channels. In (I)-(IV), CP1 and CP2 originally belonging to (B) in Fig. 8 are merged into (D), while (C) remains unchanged. In (D), CP2 can be generated by Pz, and the signal data of CP2 can be inverted to obtain CP1.
IV-E Optimized Generation Paths on Positive Emotion Data
In Fig. 9, (VI)–(IX) are the positive emotion generation paradigms. It’s also noteworthy that (VII)-(IX) exhibit similar generation paradigms. In (A) of (VI)-(IX), Fp2 serves predominantly as the reference channel, while F7 operates as the reference channel in a one-way generation paradigm alone, contrasting with the structure of (A) in section IV-D, but the paradigms in other sections align consistently with section IV-D.
TABLE II: The results of competing models for categorizing the across-subject and intra-subject nine emotions on the original dataset as well as the generated data, Where Raw_data (RAD) denotes the original dataset, and Generated_division (GD) denotes data generated based on the generative paths (depicted in Fig. 8 and 9) of neutral as well as both negative and positive sentiment generation paths. CL denotes the classification task with cross or intra subject level.
Model CL Data ACC AUC SEN SPE
SVM[51] Cross RAD 0.430$\pm$0.051 0.675$\pm$0.029 0.421$\pm$0.051 0.424$\pm$0.050
GD 0.843$\pm$0.028 0.886$\pm$0.087 0.841$\pm$0.026 0.845$\pm$0.040
Intra RAD 0.700$\pm$0.030 0.828$\pm$0.017 0.695$\pm$0.031 0.698$\pm$0.031
GD 0.940$\pm$0.052 0.966$\pm$0.029 0.939$\pm$0.051 0.940$\pm$0.050
RandomForest[52] Cross RAD 0.381$\pm$0.049 0.649$\pm$0.028 0.375$\pm$0.050 0.370$\pm$0.057
GD 0.947$\pm$0.038 0.970$\pm$0.020 0.947$\pm$0.035 0.948$\pm$0.033
Intra RAD 0.968$\pm$0.061 0.982$\pm$0.035 0.968$\pm$0.062 0.968$\pm$0.060
GD 0.973$\pm$0.056 0.985$\pm$0.032 0.972$\pm$0.056 0.973$\pm$0.055
Adaboost[53] Cross RAD 0.346$\pm$0.058 0.631$\pm$0.032 0.343$\pm$0.056 0.331$\pm$0.059
GD 0.450$\pm$0.021 0.631$\pm$0.031 0.432$\pm$0.018 0.354$\pm$0.059
Intra RAD 0.473$\pm$0.018 0.702$\pm$0.010 0.470$\pm$0.018 0.463$\pm$0.018
GD 0.460$\pm$0.023 0.624$\pm$0.037 0.441$\pm$0.024 0.396$\pm$0.114
Xgboost[54] Cross RAD 0.374$\pm$0.046 0.646$\pm$0.026 0.371$\pm$0.047 0.363$\pm$0.052
GD 0.913$\pm$0.022 0.950$\pm$0.013 0.911$\pm$0.023 0.914$\pm$0.021
Intra RAD 0.806$\pm$0.042 0.890$\pm$0.024 0.804$\pm$0.043 0.807$\pm$0.042
GD 0.944$\pm$0.067 0.968$\pm$0.038 0.943$\pm$0.068 0.946$\pm$0.065
Naive Bayes Cross RAD 0.313$\pm$0.033 0.613$\pm$0.019 0.311$\pm$0.034 0.358$\pm$0.053
GD 0.906$\pm$0.027 0.947$\pm$0.015 0.906$\pm$0.027 0.910$\pm$0.023
Intra RAD 0.364$\pm$0.014 0.642$\pm$0.008 0.363$\pm$0.014 0.438$\pm$0.025
GD 0.911$\pm$0.050 0.950$\pm$0.028 0.911$\pm$0.049 0.914$\pm$0.048
ResNet[55] Cross RAD 0.577$\pm$0.167 0.890$\pm$0.104 0.576$\pm$0.166 0.614$\pm$0.183
GD 0.951$\pm$0.053 0.996$\pm$0.006 0.950$\pm$0.056 0.963$\pm$0.040
Intra RAD 0.844$\pm$0.223 0.953$\pm$0.089 0.843$\pm$0.222 0.850$\pm$0.212
GD 0.965$\pm$0.015 0.998$\pm$0.003 0.963$\pm$0.016 0.967$\pm$0.017
EfficientNet[56] Cross RAD 0.601$\pm$0.111 0.923$\pm$0.071 0.607$\pm$0.112 0.664$\pm$0.133
GD 0.953$\pm$0.069 0.993$\pm$0.013 0.952$\pm$0.070 0.963$\pm$0.054
Intra RAD 0.812$\pm$0.183 0.960$\pm$0.072 0.818$\pm$0.112 0.842$\pm$0.153
GD 0.966$\pm$0.013 0.994$\pm$0.006 0.964$\pm$0.013 0.967$\pm$0.015
Figure 10: Visualization of spectral clustering results using t-SNE based on the raw data RAD and generated data GD.
TABLE III: The list of the combinations No.2-17 is presented in the form of ”A-B-C”, where A represents a negative emotion, B represents the neutral emotion, and C represents a positive emotion.
NO. Combination NO. Combination
1 Raw_Data (RAD) 10 Fear-Neutral-Amusement
2 Anger-Neutral-Amusement 11 Fear-Neutral-Inspiration
3 Anger-Neutral-Inspiration 12 Fear-Neutral-Joy
4 Anger-Neutral-Joy 13 Fear-Neutral-Tenderness
5 Anger-Neutral-Tenderness 14 Sadness-Neutral-Amusement
6 Disgust-Neutral-Amusement 15 Sadness-Neutral-Inspiration
7 Disgust-Neutral-Inspiration 16 Sadness-Neutral-Joy
8 Disgust-Neutral-Joy 17 Sadness-Neutral-Tenderness
9 Disgust-Neutral-Tenderness 18 Generated_division (GD)
Figure 11: The classification results using SVM with DE features to perform cross-subject recognition task of nine-category emotions on combination (1) RAD, combination (2)-(17) and combination (18) GD corresponding to Table III No.1-18 respectively.
Figure 12: The classification results using SVM with DE features to perform intra-subject recognition task of nine-category emotions on combination (1) RAD, combination (2)-(17) and combination (18) GD corresponding to Table III No.1-18 respectively.
TABLE IV: The results of competing models for categorizing the across-subject and intra-subjetc nine emotions on the original dataset as well as the generated data based on the combination No.2 of Table III, where Raw_data (RAD) denotes the original dataset, and ${\textbf{GD}_{\alpha 1}}$ denotes data generated based on the generative path (Fig. 8) of neutral, as well as negative and positive sentiment generation paths using (I) and (VI) in Fig. 9 respectively. CL denotes the classification task with cross or intra subject level.
Model CL Data ACC AUC SEN SPE
SVM[51] Cross RAD 0.430$\pm$0.051 0.675$\pm$0.029 0.421$\pm$0.051 0.424$\pm$0.050
${\textbf{GD}_{\alpha 1}}$ 0.678$\pm$0.026 0.817$\pm$0.014 0.674$\pm$0.025 0.681$\pm$0.029
Intra RAD 0.700$\pm$0.030 0.828$\pm$0.017 0.695$\pm$0.031 0.698$\pm$0.031
${\textbf{GD}_{\alpha 1}}$ 0.940$\pm$0.052 0.966$\pm$0.029 0.939$\pm$0.051 0.940$\pm$0.050
RandomForest[52] Cross RAD 0.381$\pm$0.049 0.649$\pm$0.028 0.375$\pm$0.050 0.370$\pm$0.057
${\textbf{GD}_{\alpha 1}}$ 0.818$\pm$0.049 0.887$\pm$0.053 0.815$\pm$0.047 0.818$\pm$0.056
Intra RAD 0.968$\pm$0.061 0.982$\pm$0.035 0.968$\pm$0.062 0.968$\pm$0.060
${\textbf{GD}_{\alpha 1}}$ 0.973$\pm$0.055 0.985$\pm$0.031 0.973$\pm$0.055 0.973$\pm$0.054
Adaboost[53] Cross RAD 0.346$\pm$0.058 0.631$\pm$0.032 0.343$\pm$0.056 0.331$\pm$0.059
${\textbf{GD}_{\alpha 1}}$ 0.430$\pm$0.008 0.499$\pm$0.001 0.413$\pm$0.008 0.330$\pm$0.018
Intra RAD 0.473$\pm$0.018 0.702$\pm$0.010 0.470$\pm$0.018 0.463$\pm$0.018
${\textbf{GD}_{\alpha 1}}$ 0.429$\pm$0.021 0.497$\pm$0.005 0.412$\pm$0.018 0.326$\pm$0.027
Xgboost[54] Cross RAD 0.374$\pm$0.046 0.646$\pm$0.026 0.371$\pm$0.047 0.363$\pm$0.052
${\textbf{GD}_{\alpha 1}}$ 0.805$\pm$0.053 0.889$\pm$0.028 0.802$\pm$0.050 0.812$\pm$0.040
Intra RAD 0.806$\pm$0.042 0.890$\pm$0.024 0.804$\pm$0.043 0.807$\pm$0.042
${\textbf{GD}_{\alpha 1}}$ 0.931$\pm$0.049 0.961$\pm$0.028 0.930$\pm$0.049 0.932$\pm$0.048
Naive Bayes Cross RAD 0.313$\pm$0.033 0.613$\pm$0.019 0.311$\pm$0.034 0.358$\pm$0.053
${\textbf{GD}_{\alpha 1}}$ 0.737$\pm$0.054 0.851$\pm$0.028 0.736$\pm$0.049 0.757$\pm$0.038
Intra RAD 0.364$\pm$0.014 0.642$\pm$0.008 0.363$\pm$0.014 0.438$\pm$0.025
${\textbf{GD}_{\alpha 1}}$ 0.765$\pm$0.023 0.867$\pm$0.013 0.763$\pm$0.022 0.770$\pm$0.021
ResNet[55] Cross RAD 0.577$\pm$0.167 0.890$\pm$0.104 0.576$\pm$0.166 0.614$\pm$0.183
${\textbf{GD}_{\alpha 1}}$ 0.680$\pm$0.121 0.951$\pm$0.039 0.668$\pm$0.126 0.707$\pm$0.131
Intra RAD 0.844$\pm$0.223 0.953$\pm$0.089 0.843$\pm$0.222 0.850$\pm$0.212
${\textbf{GD}_{\alpha 1}}$ 0.862$\pm$0.165 0.978$\pm$0.044 0.864$\pm$0.164 0.867$\pm$0.161
EfficientNet[56] Cross RAD 0.601$\pm$0.111 0.923$\pm$0.071 0.607$\pm$0.112 0.664$\pm$0.133
${\textbf{GD}_{\alpha 1}}$ 0.708$\pm$0.104 0.961$\pm$0.029 0.697$\pm$0.108 0.723$\pm$0.113
Intra RAD 0.812$\pm$0.183 0.960$\pm$0.072 0.818$\pm$0.112 0.842$\pm$0.153
${\textbf{GD}_{\alpha 1}}$ 0.834$\pm$0.160 0.970$\pm$0.041 0.821$\pm$0.152 0.837$\pm$0.151
IV-F Results of Emotion Recognition
In addition, we implement signal generation for each emotion according to the generation path in Fig. 8 and Fig. 9, and use SVM [51] with differential entropy (DE) features [57] to perform recognition task of nine-category emotions on the original dataset and the generated data to verify whether the generated data conforms to the data distribution of the original dataset. Besides, other classical machine learning methods[52], [53], [54] combing DE features are also used for both cross-subject emotion recognition and intra-subject emotion recognition. In intra-subject emotion recognition, each video clip’s EEG data is split 90% for training and 10% for testing, while in cross-subject emotion recognition, subjects are divided into 10 folds with a nine-fold training and one-fold testing split, repeated 10 times with classification performances averaged across 10 folds. Better yet, as illustrated in Tables II, the data generated based on the signal generation paths for the nine emotions as described in Fig. 8 and Fig. 9 exhibits remarkable efficacy by five classical machine learning methods. Furthermore, for algorithm developers, they are more inclined to design deep learning modules to compete [58], [59], [60], [61], [62], [63]. For this reason, we additionally use ResNet [55] and EfficientNet [56], without DE features to perform the nine-category emotion recognition task. The results are also shown in Table II, with remarkable improvement. These mean that our model can improve the quality of the dataset and greatly reduce data noise caused by some unexpected events during the collection process, and indicate that the EEG signals data produced by our generation path across nine emotions not only aligns with the actual data distribution but also demonstrates that, through the modeling of Hypotheses 1-3 (III-B), these paths can effectively integrate emotions with corresponding EEG signal areas for latent alignment. Moreover, for downstream tasks, the performance of simply using traditional algorithms with feature engineering to recognize generated signals can match the recognition capabilities of complex neural network models, which suggests that the generated data will facilitate the development of downstream identification models and other applications. We also use spectral clustering [64] to cluster the original data and generated data. Subsequently, we visualize the results using t-SNE [65] as shown in Fig. 10, which further proves the efficacy of our generation model.
Furthermore, we also paired the four signal generation paradigms of negative emotions with the four signal generation paradigms of positive emotions, that is, any paradigm of negative emotions was used to generate signals of other negative emotions, and the same operation was performed for positive emotions, while the signal generation method of neutral emotion remained unchanged, for a total of 16 combinations, as shown in Table III No.2-17. We adopt the same scheme as in Table II to perform recognition tasks of nine-category emotions on the generated data. The results for combination No.2 (Anger-Neutral-Amusement) are presented in Table IV, while the results for other categories are provided in the supplementary material. At the same time, Fig. 11 and 12 successively show the performance of cross-subject and intra-subject recognition task of nine-category emotions based on combination (1) raw data (RAD), combination (2)-(17) and combination (18) Generated_division (GD) responding to Table III No.1-18 respectively, using SVM with DE features. Base on these results, we can conclude that under the condition of using the same recognition model, GD is much more effective than RAD.